OpenStudy (anonymous):

What is the solution to y = 3x + 2 for the replacement set x = {1.1}?

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

so multiply 3 x 1.1

OpenStudy (anonymous):

3.3

Mehek (mehek14):

3.3 + 2

OpenStudy (anonymous):

5.3

Mehek (mehek14):

that would be the solution

OpenStudy (anonymous):

wait one sec

Mehek (mehek14):

ok

OpenStudy (anonymous):

this is what i have to put the answer into. how?

Mehek (mehek14):

oh i get it ok the already give you x which is 1.1 so put that in the first box

Mehek (mehek14):

they*

Mehek (mehek14):

now for the second box, you would put 5.3

Mehek (mehek14):

you're just making it into an ordered pair

OpenStudy (anonymous):

so it would look like?

Mehek (mehek14):

1.1, 5.3

OpenStudy (anonymous):

yes no maybe so?

Mehek (mehek14):

yes

Mehek (mehek14):

correct

OpenStudy (anonymous):

thank u so much can u help me with some more?

OpenStudy (anonymous):

What is the solution to 2.3a = 9.2?

Mehek (mehek14):

divide 9.2/2.3

OpenStudy (anonymous):

4

Mehek (mehek14):

a = 4

OpenStudy (anonymous):

thank u

Mehek (mehek14):

yw

OpenStudy (anonymous):

What is the solution to y + 7 = 15?

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

last one

Mehek (mehek14):

15 - 7

OpenStudy (anonymous):

8

Mehek (mehek14):

y = 8

OpenStudy (anonymous):

thank u so much

Mehek (mehek14):

ur very welcome and ty for the fan testimonial

Mehek (mehek14):

\(\Huge \ddot \smile\)

OpenStudy (anonymous):

no problamo

OpenStudy (anonymous):

What is the solution to y = 3x + 2 for the replacement set x = {1.1}? 1.1 , 5.3 What is the solution to 2.1m = 6.3? m = 3 What is the solution to y + 7 = 15? y = 8 Pretty sure my answers are correct. I made 100% on the test. I know that I'm like, 27 days late but it's just for others looking for the correct answers. :D